# FSAG: Enhancing Human-to-Dexterous-Hand Finger-Specific Affordance Grounding via Diffusion Models

### v1

@@ Abstract
Dexterous grasp synthesis remains a central challenge: the high dimensionality and kinematic diversity of multi-fingered hands prevent direct transfer of algorithms developed for parallel-jaw grippers. Existing approaches typically depend on large, hardware-specific grasp datasets collected in simulation or through costly real-world trials, hindering scalability as new dexterous hand designs emerge. To this end, we propose a data-efficient framework, which is designed to bypass robot grasp data collection by exploiting the rich, object-centric semantic priors latent in pretrained generative diffusion models. Temporally aligned and fine-grained grasp affordances are extracted from raw human video demonstrations and fused with 3D scene geometry from depth images to infer semantically grounded contact targets. A kinematics-aware retargeting module then maps these affordance representations to diverse dexterous hands without per-hand retraining. The resulting system produces stable, functionally appropriate multi-contact grasps that remain reliably successful across common objects and tools, while exhibiting strong generalization across previously unseen object instances within a category, pose variations, and multiple hand embodiments. This work (i) introduces a semantic affordance extraction pipeline leveraging vision–language generative priors for dexterous grasping, (ii) demonstrates cross-hand generalization without constructing hardware-specific grasp datasets, and (iii) establishes that a single depth modality suffices for high-performance grasp synthesis when coupled with foundation-model semantics. Our results highlight a path toward scalable, hardware-agnostic dexterous manipulation driven by human demonstrations and pretrained generative models.
@@ I Introduction
Building on this insight, we propose to infer a Finger-Specific Affordance Field (FSAF) that unifies the “how” and “where”: a dense, object-conditioned mapping that assigns, for each prospective surface location, structured likelihoods and role descriptors for individual fingers, jointly capturing feasible contact placement, approach direction hypotheses, and inter-finger relational compatibility. Specifically, we repurpose a frozen Stable Diffusion[29] U-Net as a semantic backbone to infer a FSAF and couple it with a kinematics-aware planner for on-table execution. Once FSAF is inferred, we compose novel grasps without additional teleoperation and transfer across heterogeneous dexterous hands via object-centric planning. This affordance-based representation enables more precise manipulation from only a small number of human demonstrations and supports deploying a single policy across multiple dexterous hands. Experiments demonstrate state-of-the-art performance on seen and unseen objects—superior affordance grounding and a 90% grasp success rate—with cross-embodiment transfer to different dexterous hands. In summary, we make the following contributions:
@@ II-A Learning from Human Demonstration Videos
Learning from human demonstration videos is central to robot learning because such data are easier to collect and encode strong task priors [22, 28]; however, a persistent embodiment gap separates human hands from robotic dexterous end-effectors. The prevailing remedy is retargeting—mapping human motions to robot kinematics [25, 43, 18]—which is effective for teleoperation and data gathering but typically still requires reinforcement learning in simulation to attain dexterous proficiency, thereby introducing a sim-to-real gap and limiting cross-embodiment generalization. Recently, a more fundamental route is to learn manipulation structure from human hand data that robots can directly consume. For parallel-jaw grippers, PointPolicy [13] uses video-derived keypoints to supervise gripper control; however, such position-only cues under-specify dexterous-hand interaction. In this work, we therefore learn finger-specific affordances from human demonstrations—object-conditioned descriptors that specify, per finger, contactable regions and approach directions over time—to represent how dexterous hands should interact with objects.
Figure 2: Pipeline overview. (1) Hyperfeature extraction: A frozen text-to-image diffusion U-Net encodes the object image with text conditioning. Multi-timestep, multi-scale activations are aggregated into hyperfeatures $A_{g}$. (2) Finger-specific affordance grounding: An FPN-style decoder maps $A_{g}$ to five per-finger likelihood maps $\hat{H}$ supervised by fingertip labels from human demonstrations. (3) Manipulation: Back-projecting GroundingDINO–SAM2 masks with FoundationStereo depth yields a partial object point cloud; peaks of $\hat{H}$ are lifted to 3D and local normals define phase-labeled waypoints (approach, closure, hold).A damped least-squares QP tracks these waypoints under joint and collision constraints to execute dexterous grasps.
@@ II-B Object-centric Representation Learning
Much of object-centric grasping is tailored to parallel-jaw grippers, representing grasps by 2D/3D locations. Such position-only cues under-specify dexterous-hand interactions [31, 7]. In addition, a subset of dexterous hand approaches relies on template matching, built around exemplars specific to the object or the task, thus having limited generalizability to novel objects/tasks [17, 10]. Contact-centric representations improve task relevance but require complex pipelines and costly data collection to obtain dexterous-hand contact regions [39, 4]. More recently, interaction-driven approaches such as CMKA [37] learn interaction regions from web imagery, yet map object keypoints to wrist, index, and little-finger anchors, effectively producing a single grasp parameterized by three points and limiting performance to on-shelf settings. These limitations motivate per-finger contact representations rather than region- or point-only cues. We learn per-finger contact cues and, via depth-and-segmentation projection, map them onto object surfaces, to further derive precise tabletop grasps.
@@ II-C Affordance Learning
The concept of affordance was popularized by psychologist James Gibson [12], which aims to highlight the actionable areas on objects, is crucial for robotic grasping and manipulation in dynamic, complex, and physical environments. Previous research has achieved remarkable progress in learning affordance knowledge from Human-Object-Interaction (HOI) images [40, 30] and videos [1, 23], language instructions [19], and 3D point clouds [6, 5]. More recently, researchers have expanded the reasoning ability of multimodal large language models into affordance learning [33, 24], achieving advanced affordance reasoning and generalization abilities. Moreover, the text-to-image diffusion model [29] shows promising affordance-aware understanding ability by generating precise hand-object images [38]. However, most approaches localize coarse object regions (e.g., boxes or masks), which is insufficient for fine-grained, finger-specific guidance required by dexterous hands—motivating representations at the per-finger contact level.
@@ III METHODS
We present a perception-to-optimization framework that learns finger-specific affordances from human demonstrations and converts them into executable dexterous grasps, as shown in Fig. 2. Semantically grounded hyperfeatures from a frozen text-to-image diffusion model [29] are decoded into five-channel per-finger likelihood maps, which guide contact selection on a reconstructed object point cloud. Local surface geometry defines waypointed approach/closure trajectories that are tracked via a damped least-squares QP under joint and collision constraints. As grasp execution is optimization-based, the pipeline transfers across dexterous hands with little or no retraining.
@@ III-B1 Feature synthesis and hyperfeature aggregation
Empirical evidence that diffusion models generate precise hand–object grasp [38] scenarios indicates that Stable Diffusion [29] has already acquired the grasp-critical feature space, integrating holistic affordance semantics with nuanced image details via its generative learning process. Inspired by prior work [38, 21], we leverage the rich semantic grounding capability of large pretrained text-to-image diffusion models to localize finger-specific contact affordances on object images. Distinct from previous approaches, we jointly harvest both visual and textual affordance semantics from Stable Diffusion. Roughly speaking, we employ a U-Net backbone comprising convolutional blocks, downsampling / upsampling paths, skip connections, and cross-attention modules. At each denoising timestep, cross-attention injects textual embeddings into the visual feature stream, encouraging intermediate activations to align with interpretable semantic concepts. Specifically, given an image–text pair $(x_{0},s)$, we encode a latent $z_{0}=E(x_{0})$ with the variational autoencoder (VAE) and follow the standard latent forward noising process
@@ III-B2 Finger-Specific Affordance Heatmap Prediction
Figure 3: Right: (1) finger-specific affordance likelihoods predicted from diffusion-derived hyperfeatures; (2) partial object point cloud reconstructed by back-projecting GroundingDINO–SAM2 segmentation with FoundationStereo depth; (3) approach vectors obtained from local surface normals around the selected contact candidates. Left: a damped least-squares QP tracks phase-labeled waypoints to execute a dexterous grasp.
@@
III-C Finger-Specific Affordance Matching for Grasp Planning
Upon obtaining the finger-specific affordances, we employ a perception-to-optimization pipeline (Fig. 3) to synthesize executable dexterous grasp motions. We deliberately eschew any learned policy so that the method can be ported across distinct dexterous hand embodiments without model modification or retraining; adopting a pure motion-planning and optimization formulation preserves cross-embodiment generality. Our perception pipeline reconstructs a partial object point cloud by first generating a dense depth map from stereo infrared (IR) using the FoundationStereo [35]. The target object mask, segmented by GroundingDINO-SAM2 [20, 27], is then back-projected using camera intrinsics to yield the final 3D point cloud. Using the finger keypoints obtained in Section III-B, we first re‑project the entire 3D point cloud onto the image plane. Then, for each finger keypoint, we use the previously described method to select the N point‑cloud projections that are closest to it in the image, forming robust fingertip contact candidates.
Approach Vector Construction: Given the predicted per-finger contact locations $c_{k}$, we estimate a consistent surface normal $\hat{n}_{k}$ by fitting a local tangent plane to the object point cloud and resolving its orientation with an object-centric rule (normal oriented toward the contact patch). We then parameterize a short, geometry-aligned 1D approach vector for each finger:
$\gamma_{k}(s)=c_{k}+s\,\hat{n}_{k},\qquad s\in\mathbb{R},$ (6)
and instantiate three phase anchors along this curve: approach ($s>0$), closure ($s\to 0^{+}$), and hold ($s<0$, introducing a small preload). For each phase $w\in\{\text{app},\text{clo},\text{hold}\}$, the Cartesian target is defined as $p_{k}^{\star}(w)=\gamma_{k}(s_{w})$. A hand-level waypoint is obtained by taking the barycentric average over all finger targets.
Task Formulation: Let $q\in\mathbb{R}^{n_{q}}$ be the full arm–hand joint vector. For fingertip pad $k\in\{1,\dots,K\}$, let $p_{k}(q)\in\mathbb{R}^{3}$ be its forward kinematics and $J_{k}(q)=\partial p_{k}/\partial q\in\mathbb{R}^{3\times n_{q}}$ its Jacobian. The goal is to drive all fingertip pads to the waypoint positions $\{p_{k}^{\star}(w)\}$. At control step $t$ with active waypoint $w_{t}$, define the per-fingertip tracking residuals $r_{k}(q_{t})=p_{k}(q_{t})-p_{k}^{\star}(w_{t})$. Additionally, we model finger underactuation via a coupling residual $e_{c}(q_{t})$ with Jacobian $J_{c}(q_{t})$, which is penalized in the objective (soft constraint):
$r_{t}=\begin{bmatrix}r_{1}(q_{t})\\[2.0pt] \vdots\\[2.0pt] r_{K}(q_{t})\\[2.0pt] r_{\mathrm{coup}}(q_{t})\end{bmatrix}=\begin{bmatrix}p_{1}(q_{t})-p_{1}^{\star}(w_{t})\\[2.0pt] \vdots\\[2.0pt] p_{K}(q_{t})-p_{K}^{\star}(w_{t})\\[2.0pt] e_{c}(q_{t})\end{bmatrix},\qquad J_{t}=\begin{bmatrix}J_{1}(q_{t})\\[2.0pt] \vdots\\[2.0pt] J_{K}(q_{t})\\[2.0pt] J_{c}(q_{t})\end{bmatrix}$,
with weight matrix $W=\mathrm{diag}(\omega_{1}I_{3},\dots,\omega_{K}I_{3},\lambda_{\mathrm{coup}}I_{m_{c}})$.
A Gauss–Newton step solves for $\Delta q_{t}$ via the convex quadratic program:
$\min_{\Delta q_{t}}$ $\tfrac{1}{2}\big\|W^{1/2}\big(J_{t}\Delta q_{t}+r_{t}\big)\big\|_{2}^{2}+\tfrac{\lambda_{\text{damp}}}{2}\|\Delta q_{t}\|_{2}^{2}$ (7)
s.t. $q_{\min}-q_{t}\leq\Delta q_{t}\leq q_{\max}-q_{t}$
$G_{\text{col}}\Delta q_{t}\leq h_{\text{col}}$
where $\lambda_{\text{damp}}$ is a Levenberg–Marquardt damping term; the first constraint enforces joint bounds and the second specifies linearized collision avoidance constraints (including table contact and self-collisions). We warm-start OSQP [32] at $q_{t}$, set $q_{t+1}=q_{t}+\Delta q_{t}^{\star}$, and recompute forward kinematics for the next step, ensuring motion continuity and avoiding commands that would place the robot in unreachable configurations.
Execution. The controller is implemented in the Mink framework [41]; solutions are computed in simulation and synchronized to the physical robot in real time.
@@ IV EXPERIMENTS
RQ2: Does the method improve the functional correctness and stability of dexterous grasps on seen and unseen objects, even for objects with weak part cues?
@@ IV-A Experimental Setup
Figure 5: Qualitative comparison of finger-specific affordance grounding. Left: Seen objects; Right: Unseen objects. Rows show ReKep, CMKA, Ours (DINO features), and Ours (Stable Diffusion). Overlays visualize five per-finger likelihood maps; colored dots indicate annotated fingertip contacts. The diffusion-based variant produces sharper, finger-disentangled hotspots aligned with functional parts and preserves localization quality on unseen tools. ReKep and CMKA each produce keypoints directly and a single-channel heatmap is generated via post-processing following CMKA, thus their overlays are rendered in a single purple color, whereas our method outputs five-channel heatmaps rendered with five distinct colors.
@@
Evaluation Metrics: For finger-specific affordance grounding, We report KLD, SIM, and NSS [2]:
@@
For Dexterous functional grasping, we follow the definitions of metrics provided in [4] and [5]: Grasping Success Rate (Suc. R.): A grasp is considered successful if the object can be lifted above 0.1 m and held stable longer than 3 seconds. Here, grasping specifically refers to grasping based on manipulation affordances.
@@ IV-B Results of Finger-Specific Affordance Grounding
Because there is no prior work that directly couples vision-language models with fine-grained, finger-specific affordance keypoints, we select two closest approaches as baselines. (1) Rekep [14] is a robotic manipulation keypoint prediction framework that extracts image features using DINO and obtains keypoints via unsupervised clustering. (2) CMKA [37] leverages exocentric (third-person) interaction images as input, applies SAM [16] for multi-scale object segmentation, and then performs clustering over the segmentation-driven feature hierarchy to learn object affordance keypoints. To ensure fairness, we train CMKA [37] using the exocentric interaction images from our dataset, keeping all other hyperparameters unchanged. For metric computation, we follow a procedure consistent with CMKA: the predicted keypoints are first rescaled to the original image resolution, then converted into heatmaps by placing Gaussian kernels (with $\sigma$ set to shorter side of image size / 64) centered at each keypoint. To validate the necessity of large-scale generative models in our framework, we conduct ablations that replace the diffusion-derived hyperfeatures with discriminative backbones CLIP [26] and DINO [3] used purely as feature extractors.
@@
Table I shows that both ReKep and CMKA fall short on finger-specific affordance grounding: ReKep’s unsupervised keypoints, learned purely from image features, lack the semantic selectivity needed to localize functionally meaningful, finger-wise contacts, while CMKA’s segmentation-driven hierarchy struggles when objects do not present distinctive, part-aligned regions. Fig. 5 (left, seen) and (right, unseen) qualitatively reveal these failure cases where on objects with weak or uniform appearance cue, e.g., the smooth, texture-poor tomato and the monochrome mouse, both baselines produce diffuse, centroid-biased blobs and scattered contact dots that either bleed across boundaries or miss functional parts. Our model simultaneously improves all three metrics—lower KLD and higher SIM/NSS—indicating more precise localization of affordance-relevant regions rather than generic centers or silhouette edges.
Holding data, labels, losses, and training schedule fixed while varying only the feature extractor, we isolate the contribution of diffusion-derived hyperfeatures. KLD increases by $1.33\times$ compared with DINO and $2.69\times$ with CLIP, while SIM drops by $14.2\%$$35.6\%$ and NSS by $9.1\%$$30.9\%$. Ours (DINO) tends to merge adjacent fingers into single peaks and produces broader activations with boundary spill-over. Besides, on unseen elongated tools (e.g., hammer), it allocates excess activation to the visually salient head rather than the grasp-functional handle. In contrast, our diffusion-based model maintains sharp, part-aligned peaks along the handle and places stabilizing secondary contacts on supporting faces. Across both seen and unseen items (e.g., screwdriver, wrench, bottle), diffusion hyperfeatures consistently suppress three dominant errors, i.e., centroid collapse, boundary spill-over, and fingertip swapping, supporting RQ1 that diffusion-derived hyperfeatures are critical for precise per-finger affordance localization.

### v2

@@ Abstract
Dexterous grasp synthesis must jointly satisfy functional intent and physical feasibility, yet existing pipelines often decouple semantic grounding from refinement, yielding unstable or non-functional contacts under object and pose variations. This challenge is exacerbated by the high dimensionality and kinematic diversity of multi-fingered hands, which makes many methods rely on large, hardware-specific grasp datasets collected in simulation or through costly real-world trials. We propose a data-efficient framework that bypasses robot grasp data collection by exploiting object-centric semantic priors in pretrained generative diffusion models. Temporally aligned and fine-grained grasp affordances are extracted from raw human video demonstrations and fused with 3D scene geometry from depth images to infer semantically grounded contact targets. We further incorporate these affordance regions into the grasp refinement objective, explicitly guiding each fingertip toward its predicted region during optimization. The resulting system produces stable, human-intuitive multi-contact grasps across common objects and tools, while exhibiting strong generalization to previously unseen object instances within a category, pose variations, and multiple hand embodiments. This work (i) introduces a semantic affordance extraction pipeline leveraging vision–language generative priors for dexterous grasping, (ii) demonstrates cross-hand generalization without constructing hardware-specific grasp datasets, and (iii) establishes that a single depth modality suffices for high-performance grasp synthesis when coupled with foundation-model semantics. Our results highlight a path toward scalable, hardware-agnostic dexterous manipulation driven by human demonstrations and pretrained generative models.
@@ I Introduction
Building on this insight, we infer a Finger-Specific Affordance Field (FSAF) that unifies the “where” and “how” of dexterous contacts: an object-conditioned field that assigns each surface location a finger-conditioned likelihood and role descriptor, capturing feasible contact placement and inter-finger compatibility. We repurpose a frozen Stable Diffusion [18] U-Net as a semantic backbone to predict FSAF from image–text inputs, and integrate it into an optimization-based grasp refinement and execution pipeline. Given the inferred FSAF, we synthesize finger-consistent grasps without additional teleoperation by warm-starting refinement toward the predicted regions, and transfer across heterogeneous dexterous hands by changing only the kinematic model (no retraining). Experiments show strong generalization to seen and unseen objects, improved affordance grounding, and high real-robot grasp success with cross-embodiment transfer.
In summary, we make the following contributions:
@@ II-A Learning from Human Demonstration Videos
Learning from human demonstration videos is central to robot learning because such data are easier to collect and encode strong task priors [14, 17]; however, a persistent embodiment gap separates human hands from robotic dexterous end-effectors. The prevailing remedy is retargeting—mapping human motions to robot kinematics [28, 12]—which is effective for teleoperation and data gathering but typically still requires reinforcement learning in simulation to attain dexterous proficiency, thereby introducing a sim-to-real gap and limiting cross-embodiment generalization. Recently, a more fundamental route is to learn manipulation structure from human hand data that robots can directly consume. For parallel-jaw grippers, PointPolicy [8] uses video-derived keypoints to supervise gripper control; however, such position-only cues under-specify dexterous-hand interaction. In this work, we therefore learn finger-specific affordances from human demonstrations—object-conditioned descriptors that specify, per finger, contactable regions to represent how dexterous hands should interact with objects.
Figure 2: Pipeline overview. (1) Hyperfeature extraction: A frozen text-to-image diffusion U-Net encodes the object image with text conditioning. Multi-timestep, multi-scale activations are aggregated into hyperfeatures $A_{g}$. (2) Finger-specific affordance grounding: An FPN-style decoder maps $A_{g}$ to five per-finger likelihood maps $\hat{H}$ supervised by fingertip labels from human demonstrations. (3) Manipulation: We reconstruct the object with SAM3+SAM3D, lift peaks of $\hat{H}$ to 3D contacts, and augment GraspQP with a finger-specific affordance term to refine stable, human-intuitive grasps.
@@ II-B Object-centric Representation Learning
Much of object-centric grasping is tailored to parallel-jaw grippers, representing grasps by 2D/3D locations. Such position-only cues under-specify dexterous-hand interactions [19, 4]. In addition, a subset of dexterous hand approaches relies on template matching, built around exemplars specific to the object or the task, thus having limited generalizability to novel objects/tasks [11]. Contact-centric representations improve task relevance but require complex pipelines and costly data collection to obtain dexterous-hand contact regions [25]. More recently, interaction-driven approaches such as CMKA [23] learn interaction regions from web imagery, yet map object keypoints to wrist, index, and little-finger anchors, effectively producing a single grasp parameterized by three points and limiting performance to on-shelf settings. These limitations motivate per-finger contact representations rather than region- or point-only cues. We learn per-finger contact cues and, via depth-and-segmentation projection, map them onto object surfaces, to further derive precise tabletop grasps.
@@ II-C Affordance Learning
However, most approaches localize coarse object regions (e.g., boxes or masks), which is insufficient for fine-grained, finger-specific guidance required by dexterous hands—motivating representations at the per-finger contact level.
@@ III METHODS
We present a perception-to-optimization pipeline that learns finger-specific affordances from human demonstrations and uses them to synthesize executable dexterous grasps (Fig. 2). We extract semantically grounded hyperfeatures using a frozen text-to-image diffusion model [18] and decode them into five per-finger likelihood maps, which provide finger-wise contact priors on the object surface. We then formulate grasp synthesis as an affordance-conditioned optimization: the per-finger priors are incorporated as a finger-wise alignment term alongside force-closure and feasibility objectives, thereby coupling grasp pose generation with the learned affordance priors.
@@ III-B1 Feature synthesis and hyperfeature aggregation
Empirical evidence that diffusion models generate precise hand–object grasp [24] scenarios indicates that Stable Diffusion [18] has already acquired the grasp-critical feature space, integrating holistic affordance semantics with nuanced image details via its generative learning process. Inspired by prior work [24, 13], we leverage the rich semantic grounding capability of large pretrained text-to-image diffusion models to localize finger-specific contact affordances on object images. Distinct from previous approaches, we jointly harvest both visual and textual affordance semantics from Stable Diffusion. Given an image–text pair $(x_{0},s)$, we encode a latent $z_{0}=E(x_{0})$ with the variational autoencoder (VAE) and follow the standard latent forward noising process
@@ III-B2 Finger-Specific Affordance Heatmap Prediction
Figure 3: Finger-Specific Affordance Field (FSAF) regions on the object induce an affordance-alignment term $E_{\mathrm{aff}}$.
@@
III-C Finger-Specific Affordance Matching for Dexterous Grasping
Object Reconstruction and Affordance Grounding.
We first segment the target object with SAM3 [2] and back-project the masked RGB-D depth into a partial object point cloud using the camera intrinsics. We then apply a reconstruction pipeline based on SAM3D [20] to obtain a posed object surface, and refine the estimated pose via ICP. The resulting object-to-camera transform is denoted as $\mathbf{T}_{\mathrm{obj}\to\mathrm{cam}}\in SE(3)$.
Using the finger-specific affordances derived in Section III-B, we re-project the object’s point cloud onto the image plane. For each predicted finger keypoint, we identify the nearest N point-cloud projections on the image, forming a set of robust fingertip contact candidates.
Grasp Refinement and Integration of Finger-Specific Affordances. Our grasp refinement module is inspired by GraspQP [29] and we introduce key innovations by refining this approach to account for finger-specific affordance cues, allowing for more precise and targeted grasp synthesis.
We minimize the following energy function to ensure the grasp is physically stable and meanwhile, aligned with the affordance regions of each finger:
$E=E_{\mathrm{fc}}+w_{\mathrm{dis}}E_{\mathrm{dis}}+w_{\mathrm{reg}}E_{\mathrm{reg}}+w_{\mathrm{aff}}E_{\mathrm{aff}},$ (6)
where $E_{\mathrm{fc}}$ is a force-closure term from GraspQP that optimizes contact force distributions to balance external wrenches under Coulomb friction. The term $E_{\mathrm{dis}}$ penalizes the distance between active contact points $\mathcal{C}^{\prime}$ and the object surface, ensuring proper contact placement, while $E_{\mathrm{reg}}$ consists of regularization terms designed to prevent hand-object penetration, joint-limit violations, and self-penetration.
We incorporate Finger-Specific Affordance Fields (FSAF) into grasp refinement as a finger-conditioned geometric prior. Affordance prior in the objective. Concretely, to instantiate the affordance regularizer in the refinement objective defined above, we add the following contact-to-region alignment term:
$E_{\mathrm{aff}}=\frac{1}{|\mathcal{C}|}\sum_{j\in\mathcal{C}}\min_{q\in S_{g(j)}}\|c_{j}-q\|_{2}^{2},$ (7)
where $\mathcal{C}$ denotes the set of contact points optimized during refinement, and $g(j)$ maps each contact $c_{j}$ to its corresponding finger-specific affordance region $S_{g(j)}$. Minimizing the full objective thus encourages contacts to stay on their predicted regions while the remaining terms enforce physical feasibility.
Initialization (warm start). We warm-start the refinement solver with an affordance-consistent contact layout: for each finger $i$, we initialize its fingertip contact $c_{i}$ near $S_{i}$ by enforcing $\mathrm{dist}(c_{i},S_{i})\leq\delta$ with $\delta=3\,\mathrm{cm}$. This narrows the search to role-consistent configurations and reduces sensitivity to local minima.
Execution. After optimizing the objective to obtain the refined grasp posture, we execute a three-phase motion plan—approach, close, and hold—to realize smooth finger trajectories and stable contact establishment. We evaluate grasp quality and task performance in Sec. IV-C.
@@ IV EXPERIMENTS
RQ2: Does the method improve grasp stability and produce more human-intuitive contact placements on seen and unseen objects, especially for categories with weak part cues?
@@ IV-A Experimental Setup
Figure 5: Qualitative comparison of finger-specific affordance grounding. Left: Seen objects; Right: Unseen objects. Rows show CMKA, Ours (DINO features), and Ours (Stable Diffusion). Overlays visualize five per-finger likelihood maps; colored dots indicate annotated fingertip contacts. The diffusion-based variant produces sharper, finger-disentangled hotspots aligned with functional parts and preserves localization quality on unseen tools. CMKA produces keypoints directly and a single-channel heatmap is generated via post-processing following CMKA, thus its overlays are rendered in a single purple color, whereas our method outputs five-channel heatmaps rendered with five distinct colors.
@@
Evaluation Metrics: For finger-specific affordance grounding, We report KLD, SIM, and NSS [1], following CMKA [23] for metric computation:
@@
For Dexterous grasping, we follow the definitions of metrics provided in [4] and [5]: Grasping Success Rate (Suc. R.): A grasp is considered successful if the object can be lifted above 0.1 m and held stable longer than 3 seconds. Here, grasping specifically refers to grasping based on manipulation affordances.
@@ IV-B Results of Finger-Specific Affordance Grounding
Figure 6: Qualitative grasp visualization. We visualize predicted grasps on five objects (bottle, banana, drill, screwdriver, and tomato). From left to right, we compare GrainGrasp, GraspQP with random initialization, GraspQP* (GraspQP initialized in the same way as ours), and our method. In the last column, the colored surface regions on the object mesh indicate our predicted finger-specific affordances.
Because there is no prior work that directly couples vision-language models with fine-grained, finger-specific affordance keypoints, we select the closest approach as the baseline. CMKA [23] leverages exocentric (third-person) interaction images as input, applies SAM [10] for multi-scale object segmentation, and then performs clustering over the segmentation-driven feature hierarchy to learn object affordance keypoints. To ensure fairness, we train CMKA [23] using the exocentric interaction images from our dataset, keeping all other hyperparameters unchanged. For metric computation, we follow a procedure consistent with CMKA: the predicted keypoints are first rescaled to the original image resolution, then converted into heatmaps by placing Gaussian kernels centered at each keypoint. To validate the necessity of large-scale generative models in our framework, We ablate feature extractors by replacing SD hyperfeatures with CLIP [16] and DINO [3]
@@
Table I shows that CMKA underperforms on finger-specific affordance grounding: its segmentation-driven hierarchy is brittle when objects lack distinctive, part-aligned visual regions. Fig. 5 (seen vs. unseen) illustrates typical failures, where predictions collapse toward object centroids or spill across boundaries, yielding contacts that miss functional parts. In contrast, our method improves all three metrics (lower KLD, higher SIM/NSS), indicating sharper, part-aligned localization rather than generic centers or silhouette edges.
Holding data, labels, losses, and the training schedule fixed while varying only the feature extractor, we isolate the contribution of diffusion hyperfeatures. Compared with DINO/CLIP, KLD increases by $1.33\times$$2.69\times$, while SIM drops by $14.2\%$$35.6\%$ and NSS by $9.1\%$$30.9\%$. Qualitatively, DINO features more often merge adjacent fingers and produce broader activations with boundary spill-over; on unseen elongated tools, they also bias toward visually salient parts rather than human-intuitive grasp regions. Diffusion hyperfeatures consistently suppress three dominant errors—centroid collapse, boundary spill-over, and fingertip swapping—supporting RQ1 that they are critical for precise per-finger affordance localization.